# Off the Planckian Locus: Using 2D Chromaticity to Improve In-Camera Color

### v2

@@ 5 Conclusion
Oracle 1D 0.00 0.01 3.92 2.14 3.19 4.69 6.63 3.01 1.77 2.64 3.78 5.17 2.99 1.72 2.59 3.71 5.16
2-CST [AdobeDNG] 1D 0.07 0.01 6.87 3.31 5.20 9.62 13.50 9.08 6.19 10.04 12.24 14.26 8.98 6.27 9.96 12.00 13.67
NN [nearestneighbors] 1D 29.53 0.01 5.42 2.83 4.38 6.63 9.81 4.26 2.27 3.42 5.05 7.37 4.22 2.37 3.55 5.17 7.65
CST-MLP 1D 1.41 0.01 4.28 2.27 3.37 4.96 7.60 3.40 1.99 2.90 4.13 5.72 3.35 1.94 2.82 4.04 5.81
NN [nearestneighbors] 2D 32.48 0.01 4.76 2.57 3.87 5.66 8.14 3.61 2.08 3.12 4.51 6.18 4.00 2.22 3.34 4.81 7.02
CST-MLP 2D 1.54 0.01 3.88 2.08 3.12 4.56 6.61 3.28 1.90 2.82 4.01 5.60 3.24 1.88 2.75 3.88 5.51
@@
Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90% Mean 25% 50% 75% 90%
Oracle 1D 0.00 0.01 1.00 0.34 0.67 1.28 2.18 1.16 0.38 0.78 1.41 2.61 1.21 0.46 0.89 1.55 2.80 0.92 0.30 0.61 1.26 2.07
2-CST [AdobeDNG] 1D 0.07 0.01 1.65 0.64 1.16 2.19 3.69 2.10 0.89 1.62 2.81 4.08 2.11 0.76 1.61 2.85 4.72 1.69 0.78 1.27 2.28 3.46
NN [nearestneighbors] 1D 3.67 0.01 1.38 0.50 0.93 1.75 2.98 1.67 0.55 1.15 2.13 3.65 1.53 0.56 1.12 1.95 3.42 1.31 0.51 0.94 1.73 2.76
CST-MLP 1D 1.41 0.01 1.29 0.52 0.93 1.57 2.68 1.47 0.53 1.05 1.84 2.99 1.45 0.60 1.11 1.86 3.10 1.13 0.48 0.81 1.41 2.33
NN [nearestneighbors] 2D 4.04 0.01 1.41 0.49 0.96 1.84 3.08 1.64 0.54 1.17 2.12 3.50 1.52 0.56 1.11 1.97 3.46 1.37 0.49 1.01 1.83 3.02
CST-MLP 2D 1.54 0.01 1.28 0.51 0.90 1.59 2.70 1.46 0.54 1.06 1.81 2.95 1.44 0.59 1.10 1.86 3.03 1.15 0.49 0.82 1.44 2.40
@@
Oracle 1D 0.00 0.01 2.29 1.21 1.94 2.88 3.99 2.60 1.41 2.21 3.08 4.74 2.67 1.41 2.26 3.58 4.61 2.08 1.20 1.79 2.54 3.57
2-CST [AdobeDNG] 1D 0.07 0.01 2.88 1.67 2.42 3.58 5.14 3.50 2.20 3.08 4.26 5.62 3.38 2.03 3.03 4.34 5.67 2.83 1.87 2.55 3.37 4.80
NN [nearestneighbors] 1D 3.67 0.01 2.57 1.45 2.19 3.21 4.64 3.07 1.79 2.66 3.83 5.46 2.91 1.58 2.46 3.89 4.98 2.42 1.49 2.12 2.95 4.06
CST-ML 1D 1.41 0.01 2.47 1.39 2.11 3.15 4.24 2.89 1.70 2.50 3.44 5.03 2.80 1.56 2.40 3.63 4.87 2.22 1.38 1.93 2.62 3.77
NN [nearestneighbors] 2D 4.04 0.01 2.62 1.44 2.23 3.28 4.76 3.06 1.70 2.53 3.70 5.60 2.91 1.62 2.47 3.84 4.89 2.47 1.46 2.16 3.04 4.31
CST-MLP 2D 1.54 0.01 2.46 1.40 2.10 3.06 4.28 2.87 1.65 2.44 3.43 5.08 2.80 1.60 2.42 3.64 4.82 2.23 1.41 1.95 2.66 3.71
@@
Figure S10: Colorimetric mapping results on ulti-illuminant data from the LSMI dataset (Galaxy camera). This figure presents an expanded qualitative evaluation, showcasing four spatially-varying multi-illuminant scenes to supplement the two shown in the main paper. Layout for each scene: The top row displays the input raw-RGB image, the white-balanced image, and the final display-referred renderings. The middle row visualizes the contrast-stretched illumination map, the first raw-RGB color chart, and its corrected renderings. The bottom row shows the per-pixel blending map, the second raw-RGB color chart, and its corresponding renderings. Methods (from left to right): 2-CST [AdobeDNG], 1D NN [nearestneighbors], our 1D CST-MLP, 2D NN [nearestneighbors], our 2D CST-MLP, and Oracle. Mean angular errors are provided at the bottom left of each rendered chart. Observations: These additional results further validate our method’s effective extension to complex scenes and confirm the consistent performance advantage of 2D chromaticity-based approaches over 1D methods.
@@
Figure S11: Colorimetric mapping results on ulti-illuminant data from the LSMI dataset (Sony camera). We provide expanded colorimetric mapping results for the Sony camera, showcasing four multi-illuminant scenes from the LSMI dataset [lsmi]. These build upon the two examples presented in the main paper. Layout for each scene: Each of the four scenes is detailed across three rows. Row 1: Original raw-RGB input, the white-balanced version, and final rendered outputs. Row 2: The illumination map, the first raw-RGB color chart, and its rendered results. Row 3: The blending map, the second raw-RGB color chart, and its rendered results. Methods (from left to right): 2-CST [AdobeDNG], 1D NN [nearestneighbors], our 1D CST-MLP, 2D NN [nearestneighbors], our 2D CST-MLP, and Oracle. Angular errors are displayed on each rendered color chart. Observations: These comprehensive visualizations reinforce the robustness of our 2D CST-MLP in handling spatially-varying illumination and its superior accuracy compared to 1D techniques across different camera sensors.
@@
Figure S12: Colorimetric mapping results on ulti-illuminant data from the LSMI dataset (Nikon camera). This figure offers further qualitative validation with four challenging multi-illuminant scenes from the LSMI dataset [lsmi], providing a more comprehensive view than the two scenes in the main paper. Layout for each scene: For each of the four scenes, the visualization is organized as follows: (top row) the input raw image, the result after white balancing, and final corrected outputs from all methods; (middle row) the illumination map and its corrected renderings of the first color chart; (bottom row) the blending map and its corresponding renderings of the second color chart. Methods (from left to right): 2-CST [AdobeDNG], 1D NN [nearestneighbors], our 1D CST-MLP, 2D NN [nearestneighbors], our 2D CST-MLP, and Oracle. Quantitative angular errors are noted on each rendered chart. Observations: As with other cameras, these extensive results for the Nikon sensor underscore the capability of our 2D CST-MLP to manage complex, mixed-lighting conditions, outperforming methods that rely on a single-dimensional illuminant representation.

### v3

@@ 5 Conclusion
Oracle 1D 0.00 0.01 3.92 2.14 3.19 6.63 3.01 1.77 2.64 5.17 2.99 1.72 2.59 5.16
2-CST [AdobeDNG] 1D 0.07 0.01 6.87 3.31 5.20 13.50 9.08 6.19 10.04 14.26 8.98 6.27 9.96 13.67
NN [nearestneighbors] 1D 29.53 0.01 5.42 2.83 4.38 9.81 4.26 2.27 3.42 7.37 4.22 2.37 3.55 7.65
CST-MLP 1D 1.41 0.01 4.28 2.27 3.37 7.60 3.40 1.99 2.90 5.72 3.35 1.94 2.82 5.81
CST-MLP (RP) 1D 2.57 0.01 4.62 2.61 3.87 7.59 3.55 2.09 3.12 5.92 4.04 2.20 3.62 6.79
NN [nearestneighbors] 2D 32.48 0.01 4.76 2.57 3.87 8.14 3.61 2.08 3.12 6.18 4.00 2.22 3.34 7.02
CST-MLP 2D 1.54 0.01 3.88 2.08 3.12 6.61 3.28 1.90 2.82 5.60 3.24 1.88 2.75 5.51
CST-MLP (RP) 2D 2.70 0.01 4.20 2.36 3.57 6.72 3.63 2.13 3.22 6.00 3.90 2.19 3.47 6.50
@@
Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90% Mean 25% 50% 90%
Oracle 1D 0.00 0.01 1.00 0.34 0.67 2.18 1.16 0.38 0.78 2.61 1.21 0.46 0.89 2.80 0.92 0.30 0.61 2.07
2-CST [AdobeDNG] 1D 0.07 0.01 1.65 0.64 1.16 3.69 2.10 0.89 1.62 4.08 2.11 0.76 1.61 4.72 1.69 0.78 1.27 3.46
NN [nearestneighbors] 1D 3.67 0.01 1.38 0.50 0.93 2.98 1.67 0.55 1.15 3.65 1.53 0.56 1.12 3.42 1.31 0.51 0.94 2.76
CST-MLP 1D 1.41 0.01 1.29 0.52 0.93 2.68 1.47 0.53 1.05 2.99 1.45 0.60 1.11 3.10 1.13 0.48 0.81 2.33
CST-MLP (RP) 1D 2.57 0.01 1.12 0.45 0.78 2.38 1.21 0.42 0.78 2.56 1.18 0.49 0.86 2.49 1.02 0.41 0.73 2.10
NN [nearestneighbors] 2D 4.04 0.01 1.41 0.49 0.96 3.08 1.64 0.54 1.17 3.50 1.52 0.56 1.11 3.46 1.37 0.49 1.01 3.02
CST-MLP 2D 1.54 0.01 1.28 0.51 0.90 2.70 1.46 0.54 1.06 2.95 1.44 0.59 1.10 3.03 1.15 0.49 0.82 2.40
CST-MLP (RP) 2D 2.70 0.01 1.11 0.44 0.77 2.39 1.20 0.39 0.79 2.56 1.16 0.47 0.82 2.47 1.01 0.44 0.74 2.04
@@
Oracle 1D 0.00 0.01 2.29 1.21 1.94 3.99 2.60 1.41 2.21 4.74 2.67 1.41 2.26 4.61 2.08 1.20 1.79 3.57
2-CST [AdobeDNG] 1D 0.07 0.01 2.88 1.67 2.42 5.14 3.50 2.20 3.08 5.62 3.38 2.03 3.03 5.67 2.83 1.87 2.55 4.80
NN [nearestneighbors] 1D 3.67 0.01 2.57 1.45 2.19 4.64 3.07 1.79 2.66 5.46 2.91 1.58 2.46 4.98 2.42 1.49 2.12 4.06
CST-MLP 1D 1.41 0.01 2.47 1.39 2.11 4.24 2.89 1.70 2.50 5.03 2.80 1.56 2.40 4.87 2.22 1.38 1.93 3.77
CST-MLP (RP) 1D 2.57 0.01 2.84 1.58 2.52 4.87 3.20 1.92 2.79 5.54 3.28 1.92 2.90 5.71 2.05 1.21 1.74 3.55
NN [nearestneighbors] 2D 4.04 0.01 2.62 1.44 2.23 4.76 3.06 1.70 2.53 5.60 2.91 1.62 2.47 4.89 2.47 1.46 2.16 4.31
CST-MLP 2D 1.54 0.01 2.46 1.40 2.10 4.28 2.87 1.65 2.44 5.08 2.80 1.60 2.42 4.82 2.23 1.41 1.95 3.71
CST-MLP (RP) 2D 2.70 0.01 3.17 1.62 2.68 5.77 2.86 1.54 2.34 5.21 3.25 1.89 2.92 5.74 2.03 1.23 1.72 3.54
@@
Figure S10: Colorimetric mapping results on multi-illuminant data from the LSMI dataset (Galaxy camera). This figure presents an expanded qualitative evaluation, showcasing four spatially-varying multi-illuminant scenes to supplement the two shown in the main paper. Layout for each scene: The top row displays the input raw-RGB image, the white-balanced image, and the final display-referred renderings. The middle row visualizes the contrast-stretched illumination map, the first raw-RGB color chart, and its corrected renderings. The bottom row shows the per-pixel blending map, the second raw-RGB color chart, and its corresponding renderings. Methods (from left to right): 2-CST [AdobeDNG], 1D NN [nearestneighbors], our 1D CST-MLP, 2D NN [nearestneighbors], our 2D CST-MLP, and Oracle. Mean angular errors are provided at the bottom left of each rendered chart. Observations: These additional results further validate our method’s effective extension to complex scenes and confirm the consistent performance advantage of 2D chromaticity-based approaches over 1D methods.
@@
Figure S11: Colorimetric mapping results on multi-illuminant data from the LSMI dataset (Sony camera). We provide expanded colorimetric mapping results for the Sony camera, showcasing four multi-illuminant scenes from the LSMI dataset [lsmi]. These build upon the two examples presented in the main paper. Layout for each scene: Each of the four scenes is detailed across three rows. Row 1: Original raw-RGB input, the white-balanced version, and final rendered outputs. Row 2: The illumination map, the first raw-RGB color chart, and its rendered results. Row 3: The blending map, the second raw-RGB color chart, and its rendered results. Methods (from left to right): 2-CST [AdobeDNG], 1D NN [nearestneighbors], our 1D CST-MLP, 2D NN [nearestneighbors], our 2D CST-MLP, and Oracle. Angular errors are displayed on each rendered color chart. Observations: These comprehensive visualizations reinforce the robustness of our 2D CST-MLP in handling spatially-varying illumination and its superior accuracy compared to 1D techniques across different camera sensors.
@@
Figure S12: Colorimetric mapping results on multi-illuminant data from the LSMI dataset (Nikon camera). This figure offers further qualitative validation with four challenging multi-illuminant scenes from the LSMI dataset [lsmi], providing a more comprehensive view than the two scenes in the main paper. Layout for each scene: For each of the four scenes, the visualization is organized as follows: (top row) the input raw image, the result after white balancing, and final corrected outputs from all methods; (middle row) the illumination map and its corrected renderings of the first color chart; (bottom row) the blending map and its corresponding renderings of the second color chart. Methods (from left to right): 2-CST [AdobeDNG], 1D NN [nearestneighbors], our 1D CST-MLP, 2D NN [nearestneighbors], our 2D CST-MLP, and Oracle. Quantitative angular errors are noted on each rendered chart. Observations: As with other cameras, these extensive results for the Nikon sensor underscore the capability of our 2D CST-MLP to manage complex, mixed-lighting conditions, outperforming methods that rely on a single-dimensional illuminant representation.